A coil inductance 8.44 mH and resistance 6 O is connected to a. 12 V battery. The current in the coil is 1.0  A at approximately the time
(a) 500 s
(b) 25 s
(c) 35 ms
(d) 1 ms

Grade:6

1 Answers

Navjyot Kalra
9 years ago

(d) KEY CONCEPT : Using I = I0

But I0 = V/R and t = L/R

? I = V/R (1 – e-Rt/L) = 12/6 [ 1 – e-6t/8.4x 10-3]

1 (given)

? t = 0.97 x 10-3 ˜ 1 ms

(D)

1. (a,b,c)

1/RC, R/L and 1/ vLC have the dimensions of frequency